OpenStudy (anonymous):

Solve this system of equations using substitution. put your answer in ordered pair form, (x,y) m+n=1000 0.05m+0.06n=57

OpenStudy (anonymous):

solve for m in top eq m=1000-n plug into bottom .05*(1000-n)+.06*n=57 n=700 so plug that into top eq m+700=1000 - 700 = 700 --------------- m=300

OpenStudy (anonymous):

How would that be in (x,y) form?

OpenStudy (anonymous):

if m is x and n is y then (300,700)